# Spectral Small-Incremental-Entangling: Breaking Quasi-Polynomial Complexity Barriers in Long-Range Interacting Systems

### v1

@@ VIII Polynomial-time simulation of 1D long-range interacting systems
  • Ground states and thermal states: We prove the existence of efficient descriptions in terms of matrix product states (MPS) and matrix product operators (MPO) with polynomial bond dimension.
@@ VIII.1 SE Strength for $k$-local Hamiltonians
The above quantity was simply treated using the inequality (11) in Ref. [44, Supplementary Lemma 1], which reduce the above inequality to the desired one (279). This completes the proof. $\square$
@@ VIII.2 MPS description of the long-range ground state
Then, we upper-bound the truncation error of the Schmidt rank in the form of
$\sum_{j>D}\left(\lambda^{(s)}_{j}\right)^{2}\leq 32D_{0}D^{-\Delta/(2\tilde{J}+\Delta)}.$ (282)
@@
where $\bar{\eta}=\eta-2$, and $c^{\ast}$ is an $\mathcal{O}(1)$ constant which depends on $g$ in (8) and $\bar{\eta}$. Note that the constant $\tilde{J}$ has been defined in Eq. (279) as an upper bound for the SE strength of $V_{AB}$.
Moreover, there exists an MPS $|{\rm M}(D)\rangle$ with the bond dimension $D$ to approximate the gapped ground state $|\Omega\rangle$ on a subset $X$ within an error of
@@ VIII.2.1 Proof of Theorem 5
where $|\tilde{\phi}_{D}\rangle$ is an arbitrary (unnormalized) quantum state which has the Schmidt rank $D$ between $A_{s}$ and $B_{s}$. Then, we utilize the result in Ref. [44, Combining Lemma 7 and Ineq. (193) in its Supplementary Note], which ensures a product state $|\phi\rangle$ that satisfies
@@
By expanding $|\phi\rangle$ such that
@@
Therefore, by letting $|\tilde{\phi}_{D}\rangle=|\phi_{D}\rangle/\langle\Omega|\phi\rangle$ in the inequality (286), we get the first main inequality (282).
@@ VIII.3 Quantum dynamics
Let $|\phi\rangle$ be an arbitrary product state. Then, we construct an MPS $|{\rm M}_{t}(D)\rangle$ with the bond dimension $D$ that approximate $e^{-iHt}|\phi\rangle$ up to an error of
@@ VIII.4.1 Review of t-DMRG
We first review the t-DMRG algorithm [59, 60, 76]. The purpose of the time-dependent DMRG (t-DMRG) algorithm is to construct a matrix product state (MPS) that approximates the time evolution
@@
    We first apply the first-order expansion
@@
$\left\|\frac{|M_{\mathcal{N}}(D)\rangle}{\||M_{\mathcal{N}}(D)\rangle\|}-e^{-iHt}|\psi_{0}\rangle\right\|$ $\leq\||M_{\mathcal{N}}(D)\rangle-e^{-iHt}|\psi_{0}\rangle\|+\frac{1-\||M_{\mathcal{N}}(D)\rangle\|}{\||M_{\mathcal{N}}(D)\rangle\|}.$ (308)
@@
$\left\|\frac{|M_{\mathcal{N}}(D)\rangle}{\||M_{\mathcal{N}}(D)\rangle\|}-e^{-iHt}|\psi_{0}\rangle\right\|\leq\frac{\epsilon(2-\epsilon)}{1-\epsilon},\quad\epsilon:=\||M_{\mathcal{N}}(D)\rangle-e^{-iHt}|\psi_{0}\rangle\|.$ (309)
@@ VIII.4.2 Error guarantee and the challenging point
Here, we consider the error accumulation in each of the processes. We prove the following lemma:
@@
Let us choose $\mathcal{N}$} such that
@@
We aim to derive an upper bound for $\epsilon_{\mathcal{N}}$. Then, we consider how the error increases with the approximation step. Given the error $\epsilon_{m}$ and the MPS $|{\rm M}_{m}(D)\rangle$, we consider $\epsilon_{m+1}$ and $|{\rm M}_{m+1}(D)\rangle$. We start from
@@
$\left(1-iH\Delta t\right)|{\rm M}_{m}(D)\rangle=\sum_{j}\lambda^{(s)}_{m,j}|\phi^{(m)}_{A_{s},j}\rangle\otimes|\phi^{(m)}_{B_{s},j}\rangle$ (320)
@@
$\left\|e^{-iH\Delta t}|{\rm M}_{m}(D)\rangle-|{\rm M}_{m+1}(D)\rangle\right\|\leq\frac{(gnt)^{2}}{{\mathcal{N}}^{2}}+\sqrt{2n}\bar{\delta}_{m}(D)$ (323)
@@
For this purpose, we here consider $e^{-iH(m+1)\Delta t}|\phi\rangle$ and consider the Schmidt decomposition between $A_{s}$ and $B_{s}$ as follows:
@@
which also implies from (325)
@@
In particular, the error grows exponentially in the number of steps $m$ (with rate $\log(1+\sqrt{2n})$), which highlights the difficulty of establishing an efficiency guarantee for t-DMRG [32].
@@ VIII.4.3 Main theorem on the precision guarantee of t-DMRG
We here resolve the problem to get a meaningful precision guarantee for the t-DMRG method: Indeed, we prove the following theorem:
@@
Let $\bar{\delta}_{m}(D)$ be the error parameter defined in Eq. (VIII.4.2). For any step $m\in\{0,\dots,\mathcal{N}-1\}$, we have
@@
Consequently, combining Lemma 19,
$\bigl{\|}e^{-iHt}|\psi_{0}\rangle-|\mathcal{M}_{\mathcal{N}}(D)\rangle\bigr{\|}\leq\frac{(gnt)^{2}}{\mathcal{N}}+\frac{\mathcal{N}}{\sqrt{D/(2n)}}e^{\tilde{J}t+(gnt)^{2}/{\mathcal{N}}}.$ (334)
Remark. From the theorem, in order to achieve the precision guarantee as $\epsilon$, we need to choose $\mathcal{N}\propto t^{2}n^{2}/\epsilon$ and $D\propto e^{\mathcal{O}(t)}n^{5}/\epsilon^{4}$, which only requires a ${\rm poly}(n/\epsilon)$ time complexity.
@@ VIII.4.4 Key Idea: Rényi-entanglement monitoring
    Importantly, the Schmidt-rank truncation itself is non-increasing for the Rényi-$1/2$ entanglement, so any entanglement growth must come from the Hamiltonian step:
    $\tilde{\zeta}_{m}\leq\zeta_{m-1}$ (339)
@@
    Given $\zeta_{m}$ for the pre-truncation state $\left(1-iH\Delta t\right)|{\rm M}_{m}(D)\rangle$ at step $m$, the tail obeys
@@
    where we use the same analyses as in (305). Thus, an upper bound on $\zeta_{m}$ (hence on $\tilde{\zeta}_{m}$) yields a quantitative bound on the truncation error at that step. Summing over steps and cuts gives the final precision guarantee for t-DMRG.
@@ VIII.4.5 Proof of Theorem 6
We use the notations in the previous section. For an arbitrary $s$, we aim to derive an upper bound for $\delta_{m,s}(D)$ in Eq. (VIII.4.2) for $\forall s$:
@@
In the following, we upper-bound $\zeta_{m}$ from above based on $\zeta_{m-1}$ (or $\tilde{\zeta}_{m}$). For this purpose, we utilize
$\zeta_{m}\leq\bar{{\mathcal{J}}}(1-iH\Delta t)\tilde{\zeta}_{m}$ (342)
Note that $\zeta_{m}$ is connected to the (1/2) Rényi entanglement of $\left(1-iH\Delta t\right)|{\rm M}_{m}(D)\rangle$. A simple estimation gives $\bar{{\mathcal{J}}}(1-iH\Delta t)\propto 1+\mathcal{O}(\left\|H\right\|\Delta t)$, which is too loose and yields exponential increase of $\tilde{\zeta}_{m}$ with respect to $m$. To refine it, we utilize the equation of
@@
$\leq 1+\tilde{J}\Delta t+(gn\Delta t)^{2}$ (346)
@@ VIII.5 Quantum Gibbs states
Our task is to implement the imaginary-time evolution to $\sum_{j=1}^{\mathcal{D}_{\Lambda}}|j_{\Lambda}\rangle\otimes|j_{\Lambda^{\prime}}\rangle$ . However, we cannot directly apply the analyses of the real-time evolution. The primary challenge occurs in the analyses of (VIII.4.5), which characterized the SE strength of $\bar{{\mathcal{J}}}(1-iH\Delta t)$. In extending it to the imaginary-time evolution, we need to consider $\bar{{\mathcal{J}}}(1-H\Delta\tau)$ with $\Delta\tau$ the decomposed unit of the total imaginary time $\beta$. We then replace Eq. (VIII.4.5) with
@@
Then, the problem is that the SE strength of $\bar{{\mathcal{J}}}(e^{-(H_{A_{s}}+H_{B_{s}})\Delta\tau})$ is not equal to $1$ unlike the real-time case; instead, we have $\bar{{\mathcal{J}}}(e^{-(H_{A_{s}}+H_{B_{s}})\Delta\beta})\propto 1+n\Delta\tau$. This also modifies the inequality (VIII.4.5) by
@@
$\kappa_{\beta}:=4\left[6+\frac{4(k+1)}{\eta-2}+2k\log_{2}(d)\right]\left\lceil\beta{\mathcal{Q}}_{0}/4\right\rceil$ (354)
@@
Remark. From the definition, we have $\kappa_{\beta}\propto\beta$, and hence in order to achieve an error $\epsilon$, we need to set the bond dimension $D$ to be as large as $(n/\epsilon)^{\mathcal{O}(\beta)}$. This is qualitatively better than the state-of-the-art result of $e^{\mathcal{O}(\beta)\log^{3}(n/\epsilon)}$ in Ref. [56]. On the other hand, it is still an open question whether one can also prove the polynomial time complexity to find such an MPS approximation. Even under the high-temperature condition (i.e., $\beta\ll 1$), the current best time complexity is given by $e^{\log^{2}(n/\epsilon)}$ [57].
@@ VIII.5.1 Proof of Theorem 7
$\left\||\rho_{\beta}\rangle-|{\rm M}_{\beta}(D)\rangle\right\|^{2}\leq 2\sum_{s=1}^{n-1}\delta_{s}^{2}(D).$ (357)
@@ VIII.5.2 Proof of Lemma 20
$V_{AB}=\sum_{j=1}^{\infty}V_{j},\quad{\rm SR}(V_{j})\leq D_{0}\quad\textrm{for}\quad\forall j,$ (372)
@@
$\sum_{r=1}^{\infty}\sum_{Z\in{\mathcal{S}}_{r}}\left\|h_{Z}\right\|\leq\sum_{r=1}^{\infty}4J_{0}r^{-\eta+1}\leq 4J_{0}\left(1+\frac{1}{\eta-2}\right)=:\tilde{g}_{0}$ (382)
@@ IX Systems with short-range interactions
Let $H_{AB}$ be a Hamiltonian with short-range interactions. For arbitrary time-evolution operator $e^{-iH_{AB}t}$, there exists an operator $U_{AB,D}$ such that $\left\|U_{AB,D}-e^{-iH_{AB}t}\right\|\leq\epsilon$, where $U_{AB,D}$ has the Schmidt rank $D$ with
@@
$\lambda_{s}\leq\exp\left(C_{1}t-\frac{C_{2}\log^{2}(s)}{t\log\log(s)}\right).$ (392)
This gives the generation of the Rényi entanglement in the form of
@@
Proof sketch. For the convenience of readers, we show the sketch of the proof in Ref. [8]. We first decompose the total system into $\Lambda=A_{1}\sqcup A_{0}\sqcup B_{0}\sqcup B_{1}$, where $|A_{0}|=|B_{0}|=\ell$. Then, one can obtain the Haah-Hastings-Kothari-Low approximation of the time evolution $e^{-iH_{AB}t}$ as follows [7] (Fig. 5):
@@
where $C,v,\mu$ is determined by the Lieb–Robinson bound. We recall that $H_{A_{0}B_{0}}$, $H_{A_{0}}$ and $H_{B_{0}}$ are the subset Hamiltonians on $A_{0}\cup B_{0}$, $A_{0}$ and $B_{0}$, respectively. Note that if the higher dimension is considered, the above upper bound depends on the surface area of $A$.

### v2

@@ VIII Simulating 1D long-range interactions with polynomial complexity
  • Ground states and thermal states: We prove the existence of efficient descriptions in terms of matrix product states (MPSs) and matrix product operators (MPOs) with polynomial bond dimension.
@@ VIII.1 SE Strength for $k$-local Hamiltonians
The above quantity was simply treated using the inequality (11) in Ref. [44, Supplementary Lemma 1], which reduces the above inequality to the desired one (280). This completes the proof. $\square$
@@ VIII.2 MPS description of the long-range ground state
Then, the truncation error of the Schmidt rank is bounded above by
$\sum_{j>D}\left(\lambda^{(s)}_{j}\right)^{2}\leq 32D_{0}D^{-\Delta/(2\tilde{J}+\Delta)},$ (283)
@@
where $\bar{\eta}=\eta-2$, and $c^{\ast}$ is an $\mathcal{O}(1)$ constant depending on $g$ in (8) and $\bar{\eta}$. Note that the constant $\tilde{J}$ has been defined in Eq. (280) as an upper bound for the SE strength of $V_{AB}$.
Moreover, there exists an MPS $|{\rm M}(D)\rangle$ with the bond dimension $D$ that approximates the gapped ground state $|\Omega\rangle$ on a subset $X$ within an error of
@@ VIII.2.1 Proof of Theorem 5
where $|\tilde{\phi}_{D}\rangle$ is an arbitrary (unnormalized) quantum state with the Schmidt rank $D$ between $A_{s}$ and $B_{s}$. Then, we utilize the result in Ref. [44, Combining Lemma 7 and Ineq. (193) in its Supplementary Note], which ensures the existence of a product state $|\phi\rangle$ satisfying
@@
By expanding $|\phi\rangle$ as
@@
Therefore, by letting $|\tilde{\phi}_{D}\rangle=|\phi_{D}\rangle/\langle\Omega|\phi\rangle$ in the inequality (287), we prove the first main inequality (283).
@@ VIII.3 Quantum dynamics
For any given $|\phi\rangle$, there always exists an MPS $|{\rm M}_{t}(D)\rangle$ with the bond dimension $D$ that approximates $e^{-iHt}|\phi\rangle$, satisfying the following error bound:
@@ VIII.4.1 Review of t-DMRG
We review the t-DMRG algorithm [59, 60, 76]. The purpose of the time-dependent DMRG (t-DMRG) algorithm is to construct a matrix product state (MPS) that approximates the time evolution
@@
    First, we apply the first-order expansion
@@
$\left\|\frac{|M_{\mathcal{N}}(D)\rangle}{\||M_{\mathcal{N}}(D)\rangle\|}-e^{-iHt}|\psi_{0}\rangle\right\|$ $\leq\||M_{\mathcal{N}}(D)\rangle-e^{-iHt}|\psi_{0}\rangle\|+\left\|\frac{|M_{\mathcal{N}}(D)\rangle}{\||M_{\mathcal{N}}(D)\rangle\|}-|M_{\mathcal{N}}(D)\rangle\right\|$
$\leq\||M_{\mathcal{N}}(D)\rangle-e^{-iHt}|\psi_{0}\rangle\|+\left(1-\||M_{\mathcal{N}}(D)\rangle\|\right).$ (309)
@@
$\left\|\frac{|M_{\mathcal{N}}(D)\rangle}{\||M_{\mathcal{N}}(D)\rangle\|}-e^{-iHt}|\psi_{0}\rangle\right\|\leq 2\epsilon,\quad\epsilon:=\||M_{\mathcal{N}}(D)\rangle-e^{-iHt}|\psi_{0}\rangle\|.$ (310)
@@ VIII.4.2 Error guarantee and the challenging point
Here, we consider the error accumulation in each of the processes. We prove the following lemma.
@@
The number of steps $\mathcal{N}$ is chosen to satisfy
@@
We aim to derive an upper bound for $\epsilon_{\mathcal{N}}$. Then, we analyze how the error increases with the approximation step. Given the error $\epsilon_{m}$ and the MPS $|{\rm M}_{m}(D)\rangle$, we consider $\epsilon_{m+1}$ and $|{\rm M}_{m+1}(D)\rangle$. We start from
@@
$\left(1-iH\Delta t\right)|{\rm M}_{m}(D)\rangle=\sum_{j}\lambda^{(s)}_{m,j}|\phi^{(m)}_{A_{s},j}\rangle\otimes|\phi^{(m)}_{B_{s},j}\rangle$. (321)
@@
$\left\|e^{-iH\Delta t}|{\rm M}_{m}(D)\rangle-|{\rm M}_{m+1}(D)\rangle\right\|\leq\frac{(gnt)^{2}}{{\mathcal{N}}^{2}}+\sqrt{2n}\bar{\delta}_{m}(D).$ (324)
@@
For this purpose, we consider $e^{-iH(m+1)\Delta t}|\phi\rangle$ and consider the Schmidt decomposition between $A_{s}$ and $B_{s}$ as follows:
@@
which, when applied to (326), implies
@@
In particular, the error grows exponentially in the number of steps $m$ (with rate $\log(1+\sqrt{2n})$), highlighting the difficulty of establishing an efficiency guarantee for t-DMRG [32].
@@ VIII.4.3 Main theorem on the precision guarantee of t-DMRG
Here, we resolve the problem of obtaining a meaningful precision guarantee for the t-DMRG method: Indeed, we prove the following theorem:
@@
Let $\bar{\delta}_{m}(D)$ be the error parameter defined in Eq. (VIII.4.2). Then, for any step $m\in\{0,\dots,\mathcal{N}-1\}$, we obtain
@@
Combining Lemma 19, the following inequality holds:
$\bigl\|e^{-iHt}|\psi_{0}\rangle-|{\rm M}_{\mathcal{N}}(D)\rangle\bigr\|\leq\frac{(gnt)^{2}}{\mathcal{N}}+\frac{\mathcal{N}}{\sqrt{D/(2n)}}e^{\tilde{J}t+(gnt)^{2}/{\mathcal{N}}}.$ (335)
Remark. According to the theorem, to achieve accuracy $\epsilon$, one may set $\mathcal{N}\propto t^{2}n^{2}/\epsilon$ and $D\propto e^{\mathcal{O}(t)}n^{5}/\epsilon^{4}$. This choice leads to a time complexity that is merely polynomial in $n/\epsilon$.
@@ VIII.4.4 Key Idea: Rényi-entanglement monitoring
    Importantly, the Schmidt-rank truncation itself is nonincreasing for Rényi-$1/2$ entanglement, so any entanglement growth must come from the Hamiltonian step:
    $\tilde{\zeta}_{m}\leq\zeta_{m-1}.$ (340)
@@
    Given $\zeta_{m}$ for the pre-truncation state $\left(1-iH\Delta t\right)|{\rm M}_{m}(D)\rangle$ at step $m$, the tail satisfies
@@
    where we use the same analysis as in (306). Thus, an upper bound on $\zeta_{m}$ (hence on $\tilde{\zeta}_{m}$) yields a quantitative bound on the truncation error at that step. Summing over steps and cuts gives the final precision guarantee for t-DMRG.
@@ VIII.4.5 Proof of Theorem 6
We use the notation introduced in the previous section. Our goal is to derive an upper bound for $\delta_{m,s}(D)$ in Eq. (VIII.4.2) that holds for any $s$:
@@
In the following, we upper-bound $\zeta_{m}$ based on $\zeta_{m-1}$ (or $\tilde{\zeta}_{m}$). For this purpose, we utilize
$\zeta_{m}\leq\bar{{\mathcal{J}}}(1-iH\Delta t)\tilde{\zeta}_{m}.$ (343)
Note that $\zeta_{m}$ is connected to the (1/2) Rényi entanglement of $\left(1-iH\Delta t\right)|{\rm M}_{m}(D)\rangle$. A simple estimate gives $\bar{{\mathcal{J}}}(1-iH\Delta t)\propto 1+\mathcal{O}(\left\|H\right\|\Delta t)$, which is too loose and leads to an exponential increase in $\tilde{\zeta}_{m}$ with $m$. To refine this, we use the following expansion:
@@
$\leq 1+\tilde{J}\Delta t+(gn\Delta t)^{2},$ (347)
@@ VIII.5 Quantum Gibbs states
Our task is to implement the imaginary-time evolution to $\sum_{j=1}^{\mathcal{D}_{\Lambda}}|j_{\Lambda}\rangle\otimes|j_{\Lambda^{\prime}}\rangle$ [8]. However, we cannot directly apply the analyses of the real-time evolution. The primary challenge occurs in the analyses of (VIII.4.5), which characterized the SE strength of $\bar{{\mathcal{J}}}(1-iH\Delta t)$. In extending it to the imaginary-time evolution, we need to consider $\bar{{\mathcal{J}}}(1-H\Delta\tau)$ with $\Delta\tau$ the decomposed unit of the total imaginary time $\beta$. We then replace Eq. (VIII.4.5) with
@@
Then, the problem is that the SE strength of $\bar{{\mathcal{J}}}(e^{-(H_{A_{s}}+H_{B_{s}})\Delta\tau})$ is not equal to $1$ unlike the real-time case; instead, we have $\bar{{\mathcal{J}}}(e^{-(H_{A_{s}}+H_{B_{s}})\Delta\tau})\propto 1+n\Delta\tau$. This also modifies the inequality (VIII.4.5) by
@@
$\kappa_{\beta}:=4\left[6+\frac{4(k+1)}{\eta-2}+2k\log_{2}(d)\right]\left\lceil\beta{\mathcal{Q}}_{0}/4\right\rceil,$ (355)
@@
Remark. From the definition, we have $\kappa_{\beta}\propto\beta$, and hence to achieve an error $\epsilon$, we need to set the bond dimension $D$ to be as large as $(n/\epsilon)^{\mathcal{O}(\beta)}$. This is qualitatively better than the state-of-the-art result of $e^{\mathcal{O}(\beta)\log^{3}(n/\epsilon)}$ in Ref. [56]. On the other hand, it is still an open question whether one can also prove the polynomial time complexity to find such an MPS approximation. Even under the high-temperature condition (i.e., $\beta\ll 1$), the current best time complexity is given by $e^{\log^{2}(n/\epsilon)}$ [57].
@@ VIII.5.1 Proof of Theorem 7
$\left\||\rho_{\beta}\rangle-|{\rm M}_{\beta}(D)\rangle\right\|^{2}\leq 2\sum_{s=1}^{n-1}\delta_{s}^{2}(D),$ (358)
@@ VIII.5.2 Proof of Lemma 20
$V_{AB}=\sum_{j=1}^{\infty}V_{j},\quad{\rm SR}(V_{j})\leq D_{0}\quad\forall j,$ (373)
@@
$\sum_{r=1}^{\infty}\sum_{Z\in{\mathcal{S}}_{r}}\left\|h_{Z}\right\|\leq\sum_{r=1}^{\infty}4J_{0}r^{-\eta+1}\leq 4J_{0}\left(1+\frac{1}{\eta-2}\right)=:\tilde{g}_{0},$ (383)
@@ IX Systems with short-range interactions
Let $H_{AB}$ be a Hamiltonian with short-range interactions. For arbitrary time‑evolution operator $e^{-iH_{AB}t}$, there exists an operator $U_{AB,D}$ such that $\left\|U_{AB,D}-e^{-iH_{AB}t}\right\|\leq\epsilon$, where $U_{AB,D}$ has the Schmidt rank $D$ with
@@
$s\leq e^{\tilde{O}\left[\sqrt{t\log(1/\lambda_{s})+t^{2}}\right]}\quad\longrightarrow\quad\lambda_{s}\leq\exp\left(C_{1}t-\frac{C_{2}\log^{2}(s)}{t\log\log(s)}\right).$ (393)
This gives the generation of the Rényi entanglement of the form
@@
Proof sketch. For the reader’s convenience, we sketch the proof in the spirit of Ref. [8]. The key ideas are: (i) decomposing the time evolution using the Haah–Hastings–Kothari–Low (HHKL) approximation [7]; and (ii) applying the Schmidt-rank bounds of Refs. [39, 40] to the Taylor expansion of the time-evolution operator restricted to the central block.
We first decompose the total system into $\Lambda=A_{1}\sqcup A_{0}\sqcup B_{0}\sqcup B_{1}$, where $|A_{0}|=|B_{0}|=\ell$. Then, one can obtain the Haah-Hastings-Kothari-Low approximation of the time evolution $e^{-iH_{AB}t}$ as follows [7] (Fig. 5):
@@
where $C,v,\mu$ is determined by the Lieb–Robinson bound. We recall that $H_{A_{0}B_{0}}$, $H_{A_{0}}$ and $H_{B_{0}}$ are the subset Hamiltonians on $A_{0}\sqcup B_{0}$, $A_{0}$ and $B_{0}$, respectively. Note that if the higher dimension is considered, the above upper bound depends on the surface area of $A$.